# vAttention: Dynamic Memory Management for Serving LLMs without PagedAttention

### v1

@@ 7.2. Portability and Performance for Decodes
Table 7. Physical memory allocation bandwidth (GB per second) for vAttention with different page sizes.
7.3. Efficacy of Physical Memory Allocation
The PyTorch caching allocator allocates memory objects (e.g., tensors) without requiring a round-trip to the kernel. In contrast, vAttention needs to invoke CUDA’s kernel driver while mapping a new physical page in a request’s KV-cache. In this section, we show that with our optimizations, vAttention can effectively meet the requirements of both the prefill and decodes phases in an LLM serving system.
Figure 11. Time to compute the prefill phase of a single prompt of 16K tokens. vAttention is as good as vLLM since it allocates memory out of the critical path. For illustration purpose, results with different page sizes show the overhead of synchronously allocating physical memory before a request’s prefill phase begins.
@@ 7.3. Efficacy of Physical Memory Allocation
Table 8. Block size (number of tokens in a newly allocated physical memory page) as a function of the page size and TP configuration. Columns on the right show how much memory can be wasted per-request in the worst-case.
Table 7 shows that even with our smallest page size of 64KB, vAttention can allocate as much as 7.6GB per second per GPU. This is more than an order of magnitude higher than the maximum memory allocation rate of 600MB per second of decodes (Figure 6). Larger page sizes and higher TP dimensions increase the memory allocation rate proportionally. This shows that vAttention is more than capable of satisfying the memory allocation bandwidth of decodes.
Further, Figure 10 shows that our optimization of overlapping memory allocation with model execution also hides the latency impact of invoking CUDA APIs. This example shows the latency of consecutive decode iterations for Llama-3-8B running with TP-1 and batch size 4. Without overlapping memory allocation with compute, new pages for the KV-cache are allocated synchronously which increase the latency of some iterations from 25 milliseconds to 41 milliseconds ($\approx$ 4 millisecond latency due to memory allocation to a single request). Note that these latency spikes occur after every 1024 iterations because we used 2MB page size for these experiments and each 2MB page contains 1024 tokens for this model configuration. When memory allocation is overlapped with the model execution of previous decode iteration, the latency effect is completely hidden.
Finally, since a prefill may require more than one page for the KV-cache, we also investigate how different memory allocation schemes impact the latency of a prefill. Figure 11 shows that allocating physical memory synchronously on-demand (when our background thread, deferred reclamation and eager allocation optimizations are all disabled) can add as much as 15% overhead with 64KB page size. Larger page sizes amortize the cost of allocation and reduce the overhead to as low as 3% (with 256KB and 2MB page sizes). vAttention further reduces the cost of allocation with deferred reclamation and eager allocation while overlapping memory allocation with compute. In most cases, these optimizations ensure that a newly arrived request can simply re-use the physical memory pages that were allocated to a previous request. Therefore, vAttention incurs negligible overhead, performing as good as vLLM for prefills.
7.4. Analysis of Memory Fragmentation
Table 8 shows the block size (defined as the minimum number of tokens in a page), and how much physical memory can be (theoretically) wasted due to over allocation in the worst-case. The worst-case occurs when a new page is allocated but remains completely unused. Further, we show each model under two TP configurations – TP-1 and TP-2 – to highlight the effect of TP dimension on block size.
vAttention allocates physical memory equivalent to the page size on each TP worker whereas the per-token physical memory requirement of a worker goes down as TP dimension increases (because KV heads get split across TP workers). Therefore, block size increases proportionally with the TP dimension. Table 8 shows that this results in the smallest block sizes of 32 (Yi-34B TP-1) to 128 (Yi-6B TP-2). In terms of the amount of physical memory, 64KB page size results in a maximum theoretical waste of only 4-15MB per request which increases to 16-60MB for 256KB page size. Overall, the important point to note is that by controlling the granularity of physical memory allocation, vAttention makes memory fragmentation insignificant. Recall that serving throughput saturates at about 200 batch size for all of our models (Figure 6). Hence, even at such large batch sizes, the maximum theoretical memory waste is at most a few GBs. Therefore, similar to vLLM, vAttention is highly effective in reducing fragmentation and allows serving using large batch sizes. However, if required, page size can be reduced further to as low as 4KB which is the minimum page size supported in almost all architectures today, including NVIDIA GPUs (B et al., 2023).
Figure 12. Illustration of code changes required to replace the prefill attention kernel of FlashAttention with that of FlashInfer. vAttention supports dynamic memory allocation transparently, enabling easy switching between kernels.
Implementation Effort
@@ 7.4. Analysis of Memory Fragmentation
The primary advantage of vAttention is portability: it enables one to seamlessly integrate new attention kernels without having to having to write a paged version of it or change the serving framework. For example, switching between the prefill or decode kernels of FlashAttention and FlashInfer requires only a few lines of code changes as shown in Figure 12. In contrast, in PagedAttention, the developers first need to write a paged attention kernel and then make significant changes in the serving framework. For example, integrating FlashInfer decode kernels in vLLM required more than 600 lines of code changes spread over 15 files (vll, 2024b, d, c). Implementing the initial paging support in FlashAttention GPU kernel also required about 280 lines of code changes (fap, 2024) and additional efforts to enable support for smaller block sizes (fas, 2024a). Given the rapid pace of innovation in LLMs, we believe it is important to reduce programming burden: production systems should be able to leverage new optimizations of the attention operator without re-writing code – similar to how optimized implementations of GEMMs are leveraged by deep learning frameworks without programmer intervention.
@@ 8. Related Work
In a recent work, GMLake (Guo et al., 2024) showed that using CUDA virtual memory support can mitigate fragmentation in DNN training jobs, increasing training batch size. In particular, GMLake uses CUDA support to coalesce multiple smaller physical memory pages into a single virtually contiguous object that can prevent out-of-memory errors for large object allocations. In contrast, vAttention is focused on avoiding fragmentation for LLM inference. Different from training, LLM inference is latency sensitive and requires smaller granularity allocations. We proposed various LLM inference specific optimizations to meet these requirements.
Optimizing LLM inference is an active area of research. Various scheduling systems have been proposed to improve different aspects of LLM serving. For example, Orca (Yu et al., 2022) and vLLM (Kwon et al., 2023) are aimed at improving serving throughput with efficient batching. Sarathi (Agrawal et al., 2023) and SplitFuse (Holmes et al., 2024) split a long prefill into multiple smaller chunks and combine decode tokens with each chunk to improve GPU compute utilization. Based on similar techniques, Sarathi-Serve (Agrawal et al., 2024) proposes stall-free batching to minimize the impact of long-running prefill iterations on decode latency. Splitwise (Patel et al., 2023), DistServe (Zhong et al., 2024) and TetriInfer (Hu et al., 2024) disaggregate the prefill and decode phases, executing them on different replicas so as to avoid interference between the prefill and decode requests. For offline inference on resource-constrained devices, FlexGen (Sheng et al., 2023) proposed a scheduling and offloading strategy to improve throughput. FastServe (Wu et al., 2023) minimizes job completion times in LLM inference using preemptive scheduling.
For all the above systems to work effectively, efficient use of GPU physical memory is essential. Since vLLM, PagedAttention has been adopted in various serving frameworks e.g., TensorRT-LLM (trt, 2023), LightLLM (lig, 2023) and kernel implementations e.g., in FlashAttention (fag, 2022) and FlashInfer (fig, 2023). In contrast, vAttention offers an alternate approach for dynamic KV-cache memory management. We show that using system support for demand paging can easily add dynamic memory management support to existing kernel implementations.
@@ 9. Conclusion
PagedAttention has emerged as the de facto standard for dynamic memory allocation in LLM inference. PagedAttention eliminates the need to reserve GPU memory ahead-of-time and therefore boosts serving throughput by fitting a larger batch size in GPU memory. While PagedAttention effectively tackles memory fragmentation, we argue that its approach of storing KV-cache in non-contiguous virtual memory introduces software complexity as well as portability and efficiency challenges. Instead, we show that using low-level system support for demand paging can avoid the pitfalls of PagedAttention. Our proposed system vAttention adds support for dynamic physical memory allocation to existing attention kernels, eliminating the need to re-write GPU code or write a memory manager in the serving framework. We show that vAttention reduces software complexity while improving portability and performance.

### v2

@@ 6.4.3. Memory allocation bandwidth
Table 7. Physical memory allocation bandwidth (GB per second) with varying allocation granularity.
6.4.4. Effect of page size.
In many applications, use of smaller pages can potentially degrade performance due to TLB thrashing (neighbourhoodawareat, ; ingens, ; hawkeye, ; architecturalsupportforgpuat, ). We find that this is not the case for LLM inference. For example, Figure 11 shows that the execution latency of an attention kernel remains largely unaffected when the page size for KV-cache is reduced to 64KB, from the original 2MB. In a separate experiment, we find that these results are also consistent with very large models e.g., Llama-3-70B and GPT-3-175B. We attribute this to the regular computation pattern of the attention operator as well as the hand-tuned implementations that explicitly try to avoid irregular memory accesses.
Further, we find that 2MB pages are good enough for many online serving scenarios where latency constraints limit the maximum batch size. However, for throughput-oriented scenarios, smaller pages may be preferred. For example, using 2MB pages could serve batch sizes of up to 187, 203 and 56 for Yi-6B, Llama-3-8B and Yi-34B on a dynamic trace (dataset: OpenChat (wang2023openchat, ), load: 7 queries per seconds). In contrast, 64KB pages helped serve batch sizes of up to 240, 258 and 68 for our models (see Figure 12).
Figure 11. Effect of page size on the performance of FlashAttention’s prefill (left) and decode (right) attention kernels (model: Llama-3-8B).
Figure 12. Maximum batch size obtained with different page sizes for a dynamic workload. Large pages limit batch size due to internal fragmentation.
Figure 13. Illustration of code changes needed to replace the prefill attention kernel of FlashAttention by FlashInfer when using vAttention for dynamic physical memory allocation.
@@
Table 8. Block size (number of tokens in a physical memory page) as a function of the page size and degree of tensor parallelism. Columns on the right show how much physical memory can be wasted per-request in the worst-case.
6.5. Programming Effort
@@
vAttention makes it feasible to replace one attention kernel with another with only a few lines of code changes in the serving framework (Figure 13). In contrast, in PagedAttention, a developer first needs to write a paged kernel and then make significant changes in the serving framework. For example, integrating FlashInfer decode kernels in vLLM required more than 600 lines of code changes in over 15 files (vllmfi1, ; vllmfi2, ; vllmfi3, ). Implementing the initial paging support in FlashAttention kernel also required $\approx 280$ lines of code changes (fapaging, ) and additional efforts to enable support for smaller block sizes (fasmall, ).
7. Discussion
Batch Size FA_Paged (Block Size = 256) FA_Paged (Block Size = 16)
Model Ratio
Yi-6B 1 0.085 0.085 1.00×1.$0$
2 0.093 0.094 1.01×1.01$\times$1.01 ×
4 0.109 0.115 1.05×1.05$\times$1.05 ×
[…]
2 0.097 0.105 1.07×1.07$\times$1.07 ×
4 0.136 0.143 1.05×1.05$\times$1.05 ×
[…]
2 0.088 0.095 1.08×1.08$\times$1.08 ×
4 0.116 0.117 1.01×1.01$\times$1.01 ×
Table 9. Impact of varying block size on the execution latency (in milliseconds) of FlashAttention decode kernel.
7.1. Further Analysis of Internal Fragmentation
vAttention uses page size as the unit of physical memory allocation as opposed to the PagedAttention’s unit of KV-cache block size. Therefore, it is natural to ask what is the relationship between page size and block size. To compute this relationship, let $S$ be the size of a single token’s per-layer K-cache (or V-cache) on a worker (see §5.1). Then, page size t translates to KV-cache block size $=t/S$. Note that $S=H\times D\times P$, where $H$ is the number of KV heads on a worker, $D$ is the dimension of each KV head and $P$ is the number of bytes based on model precision (e.g., P=2 for FP16/BF16). Therefore, block size = $t/(H\times D\times P)$.
Table 8 shows the block size for different model configurations and physical memory allocation sizes; we show each model under two TP configurations – TP-1 and TP-2 – to highlight the effect of TP dimension on block size. In addition, the table also shows how much physical memory can be (theoretically) wasted by vAttention due to internal fragmentation in the worst-case. The worst-case occurs when a new page is allocated but remains completely unused.
Context Length Batch Size FI_NonPaged FI_Paged Ratio
8K 4 0.902 0.071 12.7×12.7$\times$12.7 ×
8K 8 0.931 0.128 7.3×7.3$\times$7.3 ×
8K 16 0.943 0.229 4.2×4.2$\times$4.2 ×
[…]
16K 8 1.821 0.233 7.8×7.8$\times$7.8 ×
16K 16 1.875 0.450 4.2×4.2$\times$4.2 ×
[…]
32K 8 3.291 0.458 7.2×7.2$\times$7.2 ×
32K 16 3.365 0.891 3.8×3.8$\times$3.8 ×
Table 10. Latency (in milliseconds) of paged and non-paged decode attention kernels in FlashInfer. Ratio denotes the latency of non-paged kernel divided by the latency of paged kernel (model: Llama-3-8B, 2 A100 GPUs).
vAttention allocates physical memory equivalent to the page size on each TP worker whereas the per-token physical memory requirement of a worker goes down as TP dimension increases (because KV heads get split across TP workers). Therefore, block size increases proportionally with the TP dimension. Table 8 shows that using 2MB pages leads to large KV-cache block sizes of 1024 (Llama-3-8B and Yi-34B, TP-1) to 4096 (Yi-6B, TP-2) which could waste significant amount of physical memory (100s of MBs, per-request). Use of smaller (e.g., 64KB) pages reduces internal fragmentation proportionally, resulting in KV-cache block size of 32 (Yi-34B and Llama-3-8B, TP-1) to 128 (Yi-6B, TP-2), which in turn limit the maximum theoretical waste of only 4-15MB physical memory, per request. This shows that controlling the granularity of physical memory allocation is essential for reducing fragmentation. If required, page size can be reduced further to as low as 4KB which is the minimum page size supported in almost all architectures today, including NVIDIA GPUs.
7.2. Effect of Block Size on FlashAttention Kernel
We find that varying the block size of KV-cache also affects the execution latency of FlashAttention’s decode kernel (in addition to vLLM wherein using a larger block size increase the latency of decode kernel by up to 3×3$\times$3 × as shown in Figure 3). For example, Table 9 shows that using a smaller block size of 16 increases the latency of FlashAttention decode kernel by up to $5\%$ for Yi-6B, up to $8\%$ for Llama-3-8B and up to $9\%$ for Yi-34B. This is likely because work distribution and shared memory usage in a paged kernel needs to explicitly account for block size. Hence, writing an implementation that is equally efficient for all block sizes could be non-trivial. In fact, it is interesting to see that vLLM decode kernel performs better with small block sizes whereas FlashAttention decode kernel performs better with large block sizes. While the impact of block size on kernel latency is much less pronounced in FlashAttention compared to vLLM, a degradation in performance is still undesirable for an operator as important as attention. vAttention is free of such implementation challenges because it does not require a block table.
7.3. FlashInfer Non-paged Decode Attention Kernel
Note that we did not include FlashInfer non-paged decode kernel in our evaluation (§6). There are two reasons for this. First, the non-paged kernel has a much higher latency than the paged counterpart of FlashInfer e.g., Table 10 shows that the latency of non-paged kernel is up to 14.6×14.6$\times$14.6 × higher than the paged kernel. Further, the latency of the paged kernel follows an expected trend i.e., increasing the context length or batch size leads to a proportional increase in the execution latency. In contrast, increasing the batch size does not affect the latency of the non-paged kernel (e.g., context lengths of 8K and 32K). This behavior of the non-paged kernel is counter-intuitive and likely due to un-optimized implementation. Note that the job of a memory allocator is to simply allocate memory – it cannot accelerate a kernel. Second, the non-paged decode kernel of FlashInfer does not support variable sequence lengths – a fundamental requirement for LLM inference.
@@ 8. Related Work
Optimizing LLM inference is an active area of research. Various techniques have been proposed to improve different aspects of LLM serving like batching (orca, ; vllmsosp, ; sarathiserve2024, ), disaggregation (patel2023splitwise, ; distserve2024, ; tetriinfer, ), scheduling (flexgen, ; fastserve, ). However, central to all these techniques is the need for efficient KV-cache memory management. Since vLLM, PagedAttention has been adopted in various serving frameworks e.g., TensorRT-LLM (trtllmgithub, ), LightLLM (lightllm:github, ), and libraries e.g., FlashAttention (fagithub, ) and FlashInfer (figithub, ). In contrast, vAttention offers an alternate – which we believe is also a more principled – approach to dynamic KV-cache memory management.
In a recent work, GMLake (gmlakeasplos24, ) showed that using CUDA virtual memory support can mitigate fragmentation in DNN training jobs, increasing training batch size. In particular, GMLake uses CUDA support to coalesce multiple smaller physical memory pages into a single virtually contiguous object that can prevent out-of-memory errors for large object allocations. In contrast, vAttention is focused on avoiding fragmentation for LLM inference. Different from training, LLM inference is latency sensitive and requires smaller granularity allocations. We proposed various LLM inference specific optimizations to meet these requirements.
Similar to our approach, PyTorch also recently added support for expandable segments (expandable-segments, ) that dynamically attach physical memory pages to a pre-reserved virtual memory buffer, if enabled. However, PyTorch uses expandable segments opportunistically, using synchronous memory allocation, and allocates physical memory pages only in multiples of 2MB granularity. In contrast, vAttention caters to the specific requirements of the KV-cache.
Some other concurrent works are also motivated to optimize attention kernels (flashattnhopper, ; mirage-tensor-compiler, ; thunderkittens, ; heavyhitters, ; flashdecoding++, ; leanattention, ; flash-attention-3, ). As one would expect, new research ideas do not start with paging support in mind. vAttention would make it easier to deploy them.
@@ 9. Conclusion
In this paper, we propose vAttention for dynamic memory management in LLM serving systems. The key highlight of vAttention is that it leverages system support for demand paging instead of implementing it in user space. We present various examples to show that the vAttention approach reduces programming burden while improving portability and performance compared to the popular PagedAttention approach that many LLM serving systems use today.